# A Lot of Talk and a Badge: An Exploratory Analysis of Personal Achievements in GitHub

### 30 Dec 2023 v2

@@ 7 Limitations
First, in this study, we analyzed the activities of a sample of over 6,000 GitHub developers retrieved from 51 organizations. Because GitHub currently hosts a steadily increasing number of OSS projects that are the result of the contributions of millions of developers,101010https://octoverse.github.com/#the-world-of-open-source we acknowledge the potential threat to external validity. However, we emphasize that our study is the first to analyze personal badges; it was initiated in a timely way as soon as the new feature was released in June 2022 and completed within six months after its release. As such, we argue that the novelty and timeliness of this work compensate for the limited number of sampled developers. Furthermore, we point out that this is a study that compares the activity and profiles of developers before and after the introduction of the features of personal badges and therefore it was possible to conduct it only because we already had the necessary data collected for the before stage as part of a previous study. In other words, without the already available data regarding the 6k developers – limited or not – this study would simply not have been possible. We also point out that the sampled developers contribute to large and active software-engineered projects that were carefully mined from GitHub following the recommendations provided in (Kalliamvakou et al., 2014). Finally, given the exploratory nature of our study, we argue that the current findings help us provide an initial understanding of both the effects of badges and the developers’ perception of the new gamification element introduced in GitHub. By making the complete replication package available, we encourage the community to further our understanding of personal badges through replications and extensions of this study.
@@
Another limitation is that in our quantitative analyses (RQ1-3) we only analyzed the activities of sampled developers in public repositories. Developers can choose (i.e., not the default setting) to also include their activities in private repositories as contributing to unlocking personal achievements. However,this is a piece of information that is impossible to retrieve and therefore represents an intrinsic limitation that our study shares with all the works that rely on data retrieved from public GitHub repositories.
Finally, GitHub provides no official information about personal badges and how to unlock them. As such, all descriptions provided in this study regarding badges have been inferred by GitHub community members and contributed to public repositories. One potential limitation affecting the correlation and regression analyses is that we hypothesized the cues that personal badges might send as signals based on the perceptions reported by survey participants in their responses. Albeit to build the hypotheses we considered only the perceived signals that were reported by two or more respondents, alternative hypotheses not included in the DiD regression analysis are still possible. For example, based on survey responses, in the study we hypothesized that displaying the Quickdraw badge might be a cue of timeliness. Instead, one might also read the very short time taken to close an issue or a pull request as a negative indication that the developer might not always perform thorough code reviews. Still, we notice that the potential ‘arbitrariness’ of the cues sent by badges is an intrinsic limitation of the current implementation of the available personal badges, which lack official documentation.
@@ 8 Conclusions
In this paper, we analyzed the introduction of a new gamification feature in GitHub, namely personal achievement badges. We studied the distribution of badges, the signals they send, and the effects of their introduction. We also collected evidence of the reaction of the developers’ community to the new feature. We concluded by providing recommendations to GitHub platform designers on how to improve personal badges as a gamification mechanism and as reliable cues of developers’ ability. As the implementation of the feature appears to be in the initial stage, we will consider furthering this study as changes to badges are hopefully introduced in GitHub.
@@ Acknowledgment
Appendix A Sampled Projects
| Repository | Main language | Age | Contributors | LOC | PRs | Stars |
| adobe/react-spectrum | TypeScript | 3 | 277 | 226.39k | 2.72k | 8.95k |
| alibaba/arthas | Java | 4 | 185 | 62.41k | 724 | 32.68k |
| android/sunflower | Kotlin | 4 | 75 | 4.60k | 612 | 16.89k |
| angular/angular | TypeScript | 8 | 1,999 | 563.73k | 24.35k | 89.00k |
| ansible/ansible | Python | 10 | 6,742 | 336.49k | 49.36k | 57.84k |
| apache/echarts | TypeScript | 9 | 248 | 469.83k | 1.67k | 55.63k |
| apple/swift | C++ | 7 | 1,270 | 2,363.82k | 49.57k | 63.41k |
| aseprite/aseprite | C++ | 8 | 84 | 217.16k | 505 | 21.88k |
| atom/atom | JavaScript | 10 | 636 | 210.33k | 5.28k | 59.48k |
| BabylonJS/Babylon.js | TypeScript | 9 | 635 | 362.53k | 11.07k | 20.94k |
| capistrano/capistrano | Ruby | 13 | 274 | 14.88k | 898 | 12.50k |
| chef/chef | Ruby | 13 | 1,012 | 225.88k | 10.14k | 7.25k |
| chromium/chromium | C++ | 4 | 2,937 | 31,152.90k | 168 | 15.10k |
| crystal-lang/crystal | Crystal | 10 | 536 | 340.58k | 6.51k | 18.58k |
| deepmind/alphafold | Python | 1 | 18 | 17.43k | 79 | 10.49k |
| digitalocean/doctl | Go | 7 | 173 | 886.25k | 931 | 3.11k |
| docker/compose | Go | 9 | 145 | 25.23k | 3.70k | 29.85k |
| elastic/elasticsearch | Java | 12 | 2,045 | 3.05M | 64.62k | 64.30k |
| elixir-lang/elixir | Elixir | 11 | 1,403 | 203.40k | 7.68k | 21.77k |
| fabric/fabric | Python | 13 | 19 | 10.30k | 549 | 14.11k |
| facebook/react | JavaScript | 9 | 1,798 | 389.95k | 13.67k | 209.74k |
| fastai/fastai | Jupyter Notebook | 5 | 238 | 15.30k | 2.18k | 24.10k |
| fastlane/fastlane | Ruby | 8 | 1,522 | 593.86k | 7.19k | 37.29k |
| flutter/flutter | Dart | 7 | 1,346 | 1.30M | 45.08k | 154.68k |
| google/guava | Java | 8 | 442 | 520.74k | 1.77k | 47.75k |
| grafana/grafana | TypeScript | 9 | 2,273 | 1.39M | 37.89k | 56.02k |
| hashicorp/terraform | Go | 8 | 2,122 | 371.94k | 13.22k | 37.86k |
| Homebrew/brew | Ruby | 6 | 1,112 | 172.47k | 10.76k | 36.41k |
| ionic-team/ionic-framework | TypeScript | 9 | 486 | 423.00k | 6.51k | 49.12k |
| JabRef/jabref | Java | 8 | 626 | 212.43k | 6.11k | 3.05k |
| jekyll/jekyll | Ruby | 14 | 1,198 | 27.82k | 4.38k | 46.89k |
| JetBrains/kotlin | Kotlin | 10 | 843 | 3.52M | 5.12k | 45.06k |
| jquery/jquery | JavaScript | 13 | 338 | 39.42k | 2.86k | 57.60k |
| laravel/laravel | PHP | 11 | 780 | 1.28k | 4.55k | 73.86k |
| microsoft/vscode | TypeScript | 7 | 2,311 | 1.49M | 22.27k | 147.75k |
| MinecraftForge/MinecraftForge | Java | 10 | 547 | 98.77k | 5.17k | 6.23k |
| github/hub | Go | 13 | 252 | 22.29k | 822 | 22.47k |
| moby/moby | Go | 9 | 2,401 | 1.55M | 23.51k | 66.22k |
| Netflix/zuul | Java | 9 | 73 | 25.66k | 1.05k | 12.72k |
| nodejs/node | JavaScript | 8 | 3,700 | 9.14M | 30.71k | 96.26k |
| npm/cli | JavaScript | 4 | 871 | 5.17M | 2.24k | 7.25k |
| ohmyzsh/ohmyzsh | Shell | 13 | 2,432 | 50.08k | 6.88k | 160.24k |
| OSGeo/gdal | C++ | 10 | 539 | 1.79M | 5.12k | 3.88k |
| php/php-src | C | 11 | 1,417 | 1.46M | 9.82k | 35.60k |
| prometheus/prometheus | Go | 10 | 940 | 316.73k | 7.02k | 48.81k |
| puppetlabs/puppet | Ruby | 12 | 815 | 315.48k | 9.07k | 6.99k |
| python/cpython | Python | 5 | 2,517 | 1.71M | 41.92k | 54.25k |
| pytorch/pytorch | Python | 6 | 4,165 | 2.36M | 69.43k | 68.34k |
| rails/rails | Ruby | 14 | 6,220 | 369.79k | 31.92k | 53.09k |
| RedHatOfficial/ocp4-vsphere-upi-autom. | Jinja | 3 | 15 | 2.46k | 54 | 117 |
| ruby/ruby | Ruby | 12 | 718 | 1.71M | 7.96k | 20.51k |
| scikit-learn/scikit-learn | Python | 12 | 2,938 | 355.16k | 15.69k | 54.93k |
| Shopify/liquid | Ruby | 14 | 184 | 14.04k | 890 | 10.31k |
| SpaceVim/SpaceVim | Lua | 6 | 352 | 1.49M | 2.48k | 19.95k |
| spring-projects/spring-boot | Java | 10 | 1,239 | 457.61k | 5.76k | 68.08k |
| tensorflow/tensorflow | C++ | 7 | 4,386 | 3.72M | 23.10k | 175.80k |
| twitter/finagle | Scala | 12 | 517 | 186.45k | 539 | 8.67k |
| Yelp/paasta | Python | 7 | 231 | 130.06k | 3.23k | 1.65k |
Appendix B Survey: How do GitHub users feel about Personal Badges?
Greetings! Thank you for taking an interest in our survey. We understand how precious your time is, and that is why we made sure this survey would only take less than 5 minutes. Do not worry, your data will be collected just for research purposes: it will be anonymously stored, and once analyzed, it will be reported in aggregated form. After that, it will be deleted. We promise!
Thanks again for your help!
Background information
Tell us something about you.
1.
How many years of experience do you have in open source software? ___________________
2.
How many OSS projects are you a maintainer of? _________________________________________
3.
How many OSS projects have you contributed to? ________________________________________
4.
How many followers do you have? ___________________________________________________________
5.
Do you have a public profile, and have you kept enabled the default feature of displaying personal achievement badges on your profile page?
Mark only one oval.
1.
Yes
2.
No
Badges in general
6.
What influenced your decision to display personal achievement badges on your profile page?
Check all that apply.
1.
They look nice
2.
They display useful information
3.
I was inspired by other GitHub users
4.
I was curious about trying out a new feature
5.
It was suggested to me by other GitHub users
6.
Other: _______________________________________________
7.
Do you consider the presence of personal achievement badges in general to be an indicator of your own coding skills?
Mark only one oval.
1.
1 Not really
2.
2
3.
3
4.
4
5.
5 Very much
Badges and collaboration
8.
When looking for someone to collaborate with, I notice personal achievement badges displayed on their profile page
Mark only one oval.
1.
1 Not really
2.
2
3.
3
4.
4
5.
5 Very much
9.
I consider the presence of personal achievement badges in general to be an indicator of a developer’s ability
Mark only one oval.
1.
1 Not really
2.
2
3.
3
4.
4
5.
5 Very much
10.
The presence of personal achievement badges influenced my decision to collaborate with someone
Mark only one oval.
1.
1 Not really
2.
2
3.
3
4.
4
5.
5 Very much
11.
Please explain your previous answer: ________________________________________________________
Specific Badges
Below are the personal achievement badges discovered and unlocked so far by the GitHub community members.
12.
What personal achievement badges (displayed above) do you find most important? And what do they tell you about a GitHub user? __________________________________________
13.
Are there any personal achievement badges that you feel are missing from those displayed above? What do you think these missing personal achievement badges should reflect? ___________________________________________________________________________________________
14.
For each badge in the above picture, report the intended ‘signal’ that you think it sends to others and, if different, the actual ‘signal’ it conveys to people who visualize it. _________________________________________________________________________________________________
Disqualification
15.
Please, tell us why you decided to keep your profile page private; or, if public, why you specifically opted out of having personal achievements badges displayed on your profile. _____________________________________________________________________________________
Thank you, you answered all the questions!
We will delete this information as soon as the study is over and will never share it with anyone.
Appendix C Boxplots
Through the boxplots below, we visually confirm the presence/absence of the parallel (or common) for applying a DiD regression, that is, in the absence of treatment (i.e., before the introduction of personal badges), the dependent variable trend would be the same in both the treatment group (i.e., the developers who will eventually unlock a given type of badge) and the control group (i.e., the developers without the badges).

### 02 Feb 2024 v3

@@ 7 Limitations
First, in this study, we analyzed the activities of a sample of more than 6,000 GitHub developers retrieved from 51 organizations. Because GitHub currently hosts a steadily increasing number of OSS projects that are the result of the contributions of millions of developers,999https://octoverse.github.com/#the-world-of-open-source we acknowledge the potential threat to external validity. However, we emphasize that our study is the first to analyze personal badges; it was initiated in a timely way as soon as the new feature was released in June 2022 and completed within six months after its release. As such, we argue that the novelty and timeliness of this work compensate for the limited number of sampled developers. Furthermore, we point out that this is a study that compares the activity and profiles of developers before and after the introduction of the features of personal badges and therefore it was possible to conduct it only because we already had the necessary data collected for the before stage as part of a previous study. In other words, without the already available data on the 6k developers – limited or not – this study would simply not have been possible. We also point out that the sampled developers contribute to large and active software-engineered projects that were carefully mined from GitHub following the recommendations provided in (Kalliamvakou et al., 2014). Finally, given the exploratory nature of our study, we argue that the current findings help us to provide an initial understanding of both the effects of badges and the developers’ perception of the new gamification element introduced in GitHub.
@@
Another limitation is that in our quantitative analyses (RQ1-3) we only analyzed the activities of the sampled developers in public repositories. Developers can choose to also include their activities in private repositories as contributing to unlocking personal achievements. However, this is a piece of information that is impossible to retrieve and therefore represents an intrinsic limitation that our study shares with all the works that rely on data retrieved from public GitHub repositories.
Finally, because GitHub provides no official information on personal badges and how to unlock them, all the descriptions provided in this study regarding badges have been inferred by community members and contributed to public repositories. One potential limitation affecting the correlation and regression analyses is that we hypothesized the cues that personal badges might send as signals based on the perceptions reported by survey participants in their responses. Although we considered only the perceived signals that were reported by two or more respondents, alternative hypotheses not included in the DiD regression analysis are still possible. Still, we notice that the potential ‘arbitrariness’ of the cues sent by badges is an intrinsic limitation of the current implementation of the available personal badges, which lack official documentation.
@@ 8 Conclusions
In this paper, we analyzed the introduction of a new gamification feature in GitHub, namely personal achievement badges. We studied the distribution of badges, the signals they send, and the effects of their introduction. We also collected evidence of the reaction of the developers’ community to the new feature. We concluded by providing recommendations to the designers of the GitHub platform on how to improve personal badges as a gamification mechanism and as reliable cues of developers’ ability. As the implementation of the feature appears to be in the initial stage, we will consider furthering this study, as changes to badges are hopefully introduced in GitHub.